# Streaming Technologies and Serialization Protocols: Empirical Performance Analysis

### v1

@@ VI Discussion
We can draw several conclusions based on the experiments presented in this work. We identify the following key points from our results:
@@ VI-A Recommendations
RPC systems are faster than messaging broker systems due to the overhead of the intermediate broker. This makes RPC highly efficient for high throughput and low latency transmission of large data, although they do not offer the same delivery guarantees as message broker systems.
We found that the choice of messaging technology has a greater impact than the encoding protocol.
Protocol-based encoding methods such as Capt’n Proto and ProtoBuf perform best for complex data that can be compressed, while MessagePack is a competitive choice for smaller or random data. Protocol-based encoding methods offer the fastest serialization and best compression, with Thrift offering the best throughput and Capn’Proto offering the best compression. Binary encoding methods offer more flexibility at the cost of slower encoding speed. Among the binary encoding methods we tested, MessagePack generally performed the best. Considering text-based protocols, JSON offered the best performance due to its lightweight markup and smaller payload size compared to YAML or Avro.
Among different messaging technologies, we generally found that Apache Thrift achieves very high throughput and low latency across various scenarios. With message broker systems, RabbitMQ generally demonstrates the best performance.
Surprisingly, we did not observe much of a difference when combining different protocol-based encoding and messaging systems. We hypothesized that ProtoBuf would be most efficient when combined with the gRPC, or that Capn’Proto’s RPC implementation would perform best with Capn’Proto. However, this appears not to be the case.
Larger batch sizes facilitate higher throughput for array datasets, as shown by our throughput and batch size experiment in Figure 13, when using either a binary or protocol-based serialization method. For text-based encoding methods, adding the required markup and lack of compression destroys any advantage of batching data for transmission.
@@ VI-B Limitations and future directions
One notable limitation is that in this study we did not investigate the potential of scaling with multiple clients. Previous research has examined this aspect for message queuing systems [13]. A future study could focus on examining the reliability of various RPC technologies based on the number of consumers.
@@ VII Conclusion
In this work, we investigated 132 combinations of different encoding methods and messaging technologies. We evaluated their performance across 11 different metrics and benchmarked each combination against 6 different datasets, ranging from toy datasets to machine learning, to scientific data from the fusion energy domain. We found that messaging technology has the biggest impact on performance, regardless of over the specific serialization method used. Protocol-based methods offer the highest throughput and lowest latency but at the expense of flexibility and robustness. Protocol encoding methods offered the best performance, at the cost of flexibility. Notably, we did not see much difference when combining different protocol-based encoding and messaging systems. Finally, we found that the batch size affects the data throughput for all binary and protocol-based encoding methods.
@@
SJ: Designed and implemented the experimental framework, shaping the research methodology and contributions to the writing and conceptualization of the paper. NC: Provided the MAST data for the study and offered expertise in the fusion domain, enhancing the scientific rigor of this empirical study and editing and refining the manuscript. SK: Provided technical supervision and introduced the core idea, building upon SK’s prior work at the University of Oxford, and contributed to the writing and editing of the paper, figures, and plots.

### v2

@@ VI-A Recommendations
We found that the choice of messaging technology has a greater impact on the performance than the serialization protocol. RPC systems outperform messaging broker systems in terms of speed, primarily due to the overhead introduced by the intermediary broker in messaging systems, which adds extra processing and communication steps. RPC systems are more efficient for high throughput and low latency transmission of large datasets. However, they lack the robust delivery guarantees provided by messaging broker systems. Apache Thrift achieves high throughput and low latency across various scenarios. Among broker-based systems, RabbitMQ generally demonstrates the best performance.
Protocol-based serialization methods, such as Capt’n Proto and ProtoBuf, deliver the best performance for compressible, complex datasets, while MessagePack is a competitive choice for smaller or more random data. Protocol-based serialization methods offer the fastest serialization and best compression, with Thrift offering the best throughput and Capn’Proto offering the best compression. Binary serialization methods offer more flexibility at the cost of slower serialization speeds. Among these, MessagePack generally performed the best. In terms of text-based protocols, JSON demonstrated the best performance due to its lightweight markup and smaller payload size compared toYAML or Avro.
We observed minimal differences when combining various protocol-based serialization and messaging systems. Although we hypothesized that ProtoBuf would be most efficient when combined with the gRPC, or that Capn’Proto’s RPC implementation would perform best with Capn’Proto, our results did not support that assumption.
For array datasets, larger batch sizes resulted in higher throughput when using either a binary or protocol-based serialization methods (Figure 13). In contrast, text-based serialization methods, due to the additional markup and lack of compression, showed no such benefit from batching.
@@ VI-B Limitations and future directions
A key limitation of this study is that we did not investigate the potential of scaling with multiple clients. Previous research has examined this aspect in the context of message queuing systems [15]. Future work could focus on examining the reliability of various RPC technologies as the number of consumers increases.
@@ VII Conclusion
In this work, we investigated 143 combinations of different serialization methods and messaging technologies, assessing their performance across 11 different metrics. Each combination was benchmarked using eight different datasets, ranging from toy datasets to machine learning data and scientific data from the fusion energy domain. We found that messaging technology has the most significant impact on performance, irrespective of the serialization method used. Protocol-based methods consistently deliver the highest throughput and lowest latency, though this comes at the cost of flexibility and robustness. We observed minimal differences when combining various protocol-based serialization methods and messaging systems. Lastly, we found that batch size affects the data throughput for all binary and protocol-based serialization methods.
@@
SJ: Designed and implemented the experimental framework, shaping the research methodology and contributions to the writing and conceptualization of the paper. NC: Provided the MAST data for the study and offered expertise in the fusion domain, enhancing the scientific rigor of this empirical study and editing and refining the manuscript. SK: Provided technical supervision - introduced the core idea and contributed to the writing and editing of the paper, figures, and plots.